OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 3, 9, … if there are 12 terms?

OpenStudy (foolaroundmath):

a = 1 r = 3 n = 12 S = a(r^(n) - 1)/(r-1) = 1*(3^12 - 1)/2 Sum = (3^12 - 1)/2